[figure: 9.7 cm]

INCOMMENSURATE AND SUPERCONDUCTING PHASES IN

AN EXACTLY SOLVABLE MODEL OF STRONGLY CORRELATED ELECTRONS

Igor N. Karnaukhov Institute of Metal Physics, Vernadsky Street 36, 03142 Kiev, Ukraine

We propose a lattice integrable model for description of the phenomenon of one-dimensional incommensurability and show that the incommensurate and superconducting phases can realize simultaneously. A modified one-dimensional Hubbard model with nearest neighbor t_{1} and next nearest neighbor hoppings t_{2} has been solved and investigated. The exact ground-state phase diagram in the u − τ − n (u = U/t_{1} , U is an on-site Coulomb interaction, τ = t_{2}/t_{1}, n is the density of electrons) plane has been calculated. One finds four phases: two of them exhibit an off-diagonal long-range order and thus are superconducting and one of them exhibits incommensurate and superconducting phases. The last phase has been obtained firstly. It is shown, that the incommensurate phase is realized for ratios of the next nearest neighbor hopping integral greater than some critical value τ_{c} = 1/4 or | t_{2} |> \frac{1}{4} | t_{1} | in a defined region of the density of single carriers. This phase has been illustrated by calculations of the asymptotics of the correlation functions. In the incommensurate phase the asymptotics of the correlation functions depends on both the density of single carriers n_{1} (or the Fermi momentum) and vector Q, here Q = arccos(−\frac{1}{4τ}).

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : Igor Karnaukhov

Address for correspondence : Max-Planck Institut für Physik Komplexer Systeme, Nöthnizer Straße 38, D-01187 Dresden, Germany (from 1 April to 30 June 2002) Institute of Metal Physics, Vernadsky Street 36, 03142 Kiev, Ukraine (from 1 March to 30 March 2002)

Email address : karnaui@yahoo.com